OpenStudy (petiteme):

A motorist drives north for 35.0 minutes at 85.0 km/hr and then stops for 15.0 minutes. He then continues north, traveling 130 km in 2.00 hr. (a)What is his total displacement (b)What is his average velocity?

OpenStudy (anonymous):

I'm assuming you want your answer in Km and Hrs?

OpenStudy (petiteme):

yup :)

OpenStudy (anonymous):

Okay so You have to convert everything into like units. \[35mins=\frac{ 35}{ 60 } hours\] or 0.58 hours. \[speed=\frac{ distance}{ time} \therefore distance/displacement =speed \times time\] \[displacement =\frac{ 35}{ 60 }hours \times 85.0Km/hr = 49.58 km\] \[total displacement= 49.58 + 130 = 179.58\] Displacement is a vector and since the motorist drives only in the north direction both displacements are positive giving you the sum

OpenStudy (petiteme):

OH I get it! :) I got confused awhile ago.

OpenStudy (petiteme):

I have a question about the time. Will I minus 15 minutes?

OpenStudy (anonymous):

Is 2 hrs the total time for the trip?

OpenStudy (petiteme):

2hr is for 130 km only :)

OpenStudy (petiteme):

thanks for the help anyway! you are so great :)